OpenStudy (anonymous):

Simplify 1+cos(x+pi/2)+cos(x-pi/2) , steps .

OpenStudy (ash2326):

1+cos(x+pi/2)+cos(x-pi/2) we know cos (-x)=cos (x) using this for the last term 1+cos(pi/2+x)+cos(pi/2-x) now cos(pi/2+x)=-sin x and cos (pi/2-x)=sin x using these in the equation 1-sin x+sin x 1

OpenStudy (ash2326):

simplified form =1

OpenStudy (anonymous):

Thanks!